# One Homography is All You Need: IMM-based Joint Homography and Multiple Object State Estimation

### v3

@@ 1 Introduction
The increasing reliability of object detectors [1, 2, 3, 4, 5, 6, 7] has encouraged research on the topic of tracking objects in the image space, i.e. where objects are tracked in pixel coordinates. Most recent methods have focused on using bounding box information with or without appearance embeddings [8, 9, 10, 11, 12]. Yet, it is possible in some cases to obtain an estimate of the location of a bounding box within the ground plane of the imaged 3D space [13]. The ground plane is a coordinate system representing a 2D approximation of the earth’s surface in front of the camera, as viewed from above. Leveraging such information may improve association performance. Occlusion and camera motion effects may be mitigated if the target dynamics within the ground plane are reliably estimated and the homographic projection of the ground plane to the image plane is decoupled from the target’s motion. This is illustrated in Figure 1, which depicts two tracks that undergo heavy occlusion in the image plane at time $t$. Although the red track is occluded by the blue track, their locations on the ground plane are still easily separable. Furthermore, camera motion does not influence target association, i.e. additional camera motion compensation is not required if an accurate projection matrix (homography) can be determined at any time step.
@@
In addition, many 3D MOT methods use regular 3D measurements [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], assume the internal camera parameters are known [29, 30] or attempt to estimate 3D coordinates based on 2D images [31]. Since IMM-JHSE requires a single initial homography estimate (which only provides a mapping to the 2D world plane), it is considered distinct from these approaches.
[…]
@@
This paper expands upon previous multiple object tracking (MOT) methods. In particular, it makes the following contributions:
@@
    Modelling limitations of a previous method [32] are addressed; see Section 4. Briefly, target motion is decoupled from the homography matrix state.
@@
    The proposed method outperforms related methods on the DanceTrack [33] and KITTI-car [34] datasets while offering competitive performance on the MOT17 [35], MOT20 [36], and KITTI-pedestrian [34] datasets.
@@ 2 Previous Work
IMM-JHSE uses only a single initial homography estimate as in [32]. It is important to note that the homography transformation only provides a plane-to-plane mapping. This differentiates it from fully-3D MOT methods which use regular 3D measurements obtained with LiDAR sensors (point clouds) [14, 15, 16, 17, 18, 19], RGB-D cameras [37, 38], 3D detections [20, 21, 22, 23, 24, 25, 26], odometry [28] and stereo or binocular cameras [27, 39, 40].
In addition, the methods in [29, 30] use a Reversible Jump Markov Chain Monte Carlo particle filter with various feature representations, including skin colour, depth images, facial detection [29] and explicit occlusion reasoning [30]. Furthermore, these methods consider different camera motion models than the one proposed here (which explicitly considers camera position and rotation) and assume that internal camera parameters are known. Finally, [30] further makes use of odometry information.
Furthermore, the method in [31] attempts to perform 3D tracking without any of these measurements. However, the authors take an iterative approach to camera motion estimation. In contrast, IMM-JHSE integrates camera motion and its dynamics into the state vector to estimate it jointly with track states. They also rely on detecting vanishing points and lines parallel to the horizon. In addition, the multi-mode filter they employ only considers the target state and needs to be reinitialised with previous measurements when a mode switch occurs. Finally, multiple hypothesis tracking (MHT) performs association, whereas IMM-JHSE uses a linear assignment algorithm.
@@ 2.2 Tracking by attention
Instead of the pre-defined anchor boxes used in several popular object detectors, Carion et al. [7] decode object positions from a set of learned position embeddings using a transformer decoder. Meinhardt et al. [41] extend this approach to MOT by using previous detections as queries in subsequent frames, resulting in the Trackformer architecture. Whereas Trackformer works on a frame-by-frame basis, Global Tracking Transformers, as proposed by Zhou et al. [42], encode detections in a sliding temporal window of frames. A transformer decoder then associates new detections with these encodings. Having access to an entire measurement sequence, the deep data association module proposed by Pinto et al. [43] eschews the transformer decoder and relies solely on an encoder to associate detections within the given sequence with one another. End-to-end attention-based methods have recently succeeded on MOT leaderboards [44, 45, 46]. However, these methods are computationally expensive, especially considering they often train their own detectors in addition to training the attention-based mechanism in an end-to-end approach.
@@ 2.3 Image-based Tracking
Tracking the motion of objects from a video has been of interest as early as 1998 when Isard and Blake [47] developed the condensation algorithm for tracking objects represented by a curve or sets of curves with learned motion models. Their method propagates conditional density distributions similar to the approach of particle filtering. However, tracking multiple objects simultaneously with their method may become infeasible, especially for complex multi-modal posterior densities where large samples or particles may be required to obtain accurate estimates. Besides, multi-object tracking requires solving the association problem – deciding which observations belong to which tracked object – which becomes increasingly difficult when the tracked objects are near one another, partially or fully occluding each other. Multiple object tracking (MOT) methods in the literature focus more on the association problem than on detection or individual target tracking.
@@
Given a set of object detections, Li et al. [48] model the association problem with a cost-flow network, with the solution obtained by a minimum-cost flow algorithm. They also explicitly model occluded observations, which is achieved by simply comparing differences in detection locations and scale to predetermined thresholds. While the approach is reported to be real-time and yields favourable results on the Context-Aware Vision using Image-based Active Recognition (CAVIAR) [49] dataset, object motion is not explicitly modelled. As a result, the method may perform poorly when object motion is highly dynamic. Berclaz et al. [50] similarly model object trajectories with a flow model consisting of a discretised 2D spatial grid (representing the image area) at each time step. They use the K-shortest path algorithm to solve their linear programming formulation of the association problem. Motion is modelled by only allowing an occupied node at the previous time step to transition to a specific neighbourhood around that node in the current time step. As with [48], this is an unsatisfactory solution in situations where objects are highly dynamic.
@@
With the increasing reliability and popularity of several object detectors [1, 2, 3, 4, 5, 6, 7], modern methods have primarily made use of bounding box information. Bewley et al. [8] emphasise the importance of detection quality and use the Hungarian algorithm to perform data association based on the intersection-over-union (IoU) of detected bounding boxes with those predicted by Kalman-filtered tracks (specifically, predicted bounding boxes are constructed from the predicted bounding box centre x- and y-coordinates, height and aspect ratio). Their method has been influential in promoting the use of Kalman filters to estimate object motion distributions and/or the Hungarian algorithm (or other linear assignment algorithms) to perform association: these have become the status quo in [51, 52, 53, 54, 55, 9, 10]. Taking the minimalist approach of the original SORT algorithm one step further, Bergmann et al. [56] rely purely on the regression head of their detector network (or “Tracktor”) to perform tracking. However, they show that performance improves with camera motion compensation and appearance-feature-based re-identification, suggesting that explicit motion models (and appearance information) can still be beneficial. Supporting this idea, Khurana et al. [52] use motion models that consider monocular depth estimates to maintain occluded tracks. SORT is improved upon in [57] by integrating appearance information, resulting in DeepSORT. They train a convolutional neural network (CNN) to discriminate between pedestrians and keep each track’s normalised appearance embeddings of the last 100 frames. Track association is performed by finding the minimum cosine distance of the current detected appearance embedding with the history of embeddings for each of the previous tracks. Additionally, they use motion information by obtaining the Mahalanobis distance between detections and predicted Kalman filter states, consisting of bounding box centre positions, aspect ratios, heights and their corresponding velocities. In SiMIlarity LEarning for Occlusion-Aware Multiple Object Tracking (SMILEtrack), Hsiang et al. [58] make use of a Siamese-based network which incorporates elements inspired by vision transformers (specifically, self-attention computed on image patches) to extract appearance features from detected bounding boxes. However, it does not achieve a convincing improvement in higher-order tracking accuracy (HOTA) [59] compared to ByteTrack, despite its superior computational speed. Leveraging multiple-camera traffic surveillance systems, another method explicitly models inter-vehicle occlusion using reconstructed-cuboid projections instead of relying on similarity learning [60]. It is noted that the use of camera motion compensation has become ubiquitous, with applications in [56, 52, 51, 55, 9, 10, 58, 61].
@@
Whereas previous methods separately perform detection and appearance embedding extraction, Wang et al. [53] design a model which performs both of these simultaneously. They report that similar results can be achieved compared to state-of-the-art SDE (separate detection and embedding) methods at reduced computational cost. Zhang et al. [54] show that such multi-task networks must be designed carefully since the tasks require different features and feature dimensions and that the anchors used for object detection can introduce ambiguity in the re-identification features. They introduce a carefully designed anchor-free model to simultaneously perform object detection and appearance embedding extraction, which outperforms the state-of-the-art methods in tracking metrics and frame rate.
@@
Shifting the focus from appearance embeddings, Zhang et al. [9] note that it is beneficial to perform association in two steps in their method called ByteTrack: the first step takes only high-confidence detections into account, while the remaining unmatched detections, as well as the lower-confidence detections, are associated in a second association step. They note that using the IoU is essential in the second association step since other features may become unreliable for low-confidence detections, which may be occluded and/or blurred. Still, re-identification or other features may be used in the first step. By altering the Kalman filter state vector, Aharon et al. [10] improve upon ByteTrack. Instead of tracking the aspect ratio of bounding boxes, they perform better by directly tracking bounding box widths and heights. Similar to [55, 54, 53], an exponential moving average mechanism is used to update the appearance states of tracklets – the features of which are obtained by a re-identification network from the FastReID library [62]. Specifically, the bags of tricks [63] stronger baseline with a ResNeSt50 [64] backbone is used. Accordingly, they name their method BoT-SORT. Although tracking width and height may improve performance, Cao et al. [11] note that previous motion models are estimation-centric, allowing for rapid error accumulation over time, especially when no observations are available, and motion is non-linear. This implies that previous approaches are more sensitive to the noise in state estimation than the noise in measurements. They propose observation-centric SORT (OC-SORT), which incorporates a re-update step when an object is re-identified, during which the Kalman filter is updated with virtual detections generated based on the last-known and currently detected bounding box state. In addition, they use a novel association criterion that considers the change in motion direction that each new detection would induce, which is termed observation-centric momentum (OCM). Finally, Yang et al. [12] buffers object bounding boxes to expand the matching space of subsequent detections in their Cascaded Buffered IoU (C-BIoU) tracker. A cascaded matching approach is used such that small buffers are used in the first association step, and larger buffers are used in the second step. Despite its simplicity, Their method performs superior to DeepSORT, SORT, ByteTrack and OC-SORT.

### v4

@@ 1 Introduction
The increasing reliability of object detectors (yolo; yolo9000; yolov3; yolov4; yolov9; Ge2021YOLOX:2021; end_end_object) has encouraged research on the topic of tracking objects in the image space, i.e. where objects are tracked in pixel coordinates. Most recent methods have focused on using bounding box information with or without appearance embeddings (Bewley2016SimpleTracking; Zhang2022ByteTrack:Box; botsort; Cao2022Observation-CentricTracking; Yang2022HardSpace). All of these methods predict the bounding box dimensions of a track in the next frame and associate measured bounding boxes with existing tracks based on these predictions. Yet, it is possible in some cases to obtain an estimate of the location of a bounding box within the ground plane of the imaged 3D space (bhitk). The ground plane is a coordinate system representing a 2D approximation of the earth’s surface in front of the camera, as viewed from above. Leveraging such information may improve association performance. Occlusion and camera motion effects may be mitigated if the target dynamics within the ground plane are reliably estimated and the homographic projection of the ground plane to the image plane is decoupled from the target’s motion. This is illustrated in Figure 1, which depicts two tracks that undergo heavy occlusion in the image plane at time $t$. Although the red track is occluded by the blue track, their locations on the ground plane are still easily separable. Furthermore, camera motion does not influence target association, i.e. additional camera motion compensation is not required if an accurate projection matrix (homography) can be determined at any time step.
Despite this, tracks may momentarily leave the ground plane in certain situations, for example, dancers jumping in the DanceTrack dataset (dance). During these moments, the homographic projection of a target’s position in the image plane to the ground plane no longer accurately represents its actual ground-plane location. A multiple object tracking (MOT) algorithm that relies on homographic projections should treat these situations with care. This paper hypothesises that such an MOT algorithm could be made robust to these situations by leveraging a combination of ground- and image-plane filters. Furthermore, it is hypothesised that dynamic process noise estimation can improve performance by helping to account for the divergence of target motion from the prescribed motion model. Similarly, dynamic measurement noise estimation could improve performance by exploiting variance in measurement precision for different targets (i.e. the variability in measurements over time could differ for different targets).
Few 2D MOT methods experiment with incorporating limited 3D information. One exception is UCMCTrack (ucmc), which uses an initial homography estimate to perform target tracking in the ground plane. Despite this novelty, UCMCTrack uses the same camera motion compensation technique that has been proven beneficial in state-of-the-art 2D MOT methods like BoT-SORT (botsort). However, as will be shown, this method does not allow for a dynamic estimate of the homography matrix, induces a forced coupling between camera motion and target state, and precludes joint target and homography state estimation. In contrast, the proposed method decouples camera motion from target state and allows for the joint estimation of target and camera dynamics.
@@
Many 3D MOT methods use regular 3D measurements (Jiang2024ATracking; Osep2018CombinedScenes; Zhang2024LEGO:Clouds; Wu2022CasA:Clouds; Nagy2024RobMOT:PointCloud; Huang2024BiTrack:Data; Reich2021MonocularTracks; Weng20203DMetrics; Wu20223DAssociation; Wang2024YouTracking; Wang2023TowardsMOT; Wu2023VirtualDetection; Wang2024MCTrack:Driving; Ninh2024CollabMOTTracking; Gao2021GlowEnvironments), assume the internal camera parameters are known (Choi2013ACamera; Wojek2013MonocularScenes) or attempt to estimate 3D coordinates based on 2D images (Wu2023HybridDevices). Since IMM-JHSE requires a single initial homography estimate (which only provides a mapping to the 2D world plane), it is considered distinct from these approaches.
ground planeOptical centreimage planet1t-1t1t-1ttttt+1t+1t+1t+1t+1t+1t+1t+1ttttt1t-1t1t-1
@@
This paper expands upon previous MOT methods. In particular, it makes the following contributions:
@@
    Modelling limitations of a previous method (ucmc) are addressed; see Section 4. Briefly, target motion is decoupled from the homography matrix state.
@@
    The proposed method outperforms related methods on the DanceTrack (dance) and KITTI-car (kitti) datasets while offering competitive performance on the MOT17 (mot17), MOT20 (mot20), and KITTI-pedestrian (kitti) datasets.
In making the above-mentioned contributions, association performance may be improved in situations where an initial homography or camera calibration estimate is available in lieu of regular 3D measurements and the required hardware. Thus, the affordability of MOT may be improved in some use cases, for example, in pedestrian and car tracking, such as with the KITTI dataset (kitti), which is important for applications like self-driving cars, or for athletic and gymnastic applications like sports player tracking, dancing or gymnastic performances (e.g. the DanceTrack dataset (dance)). IMM-JHSE may be combined with homography estimation techniques, such as in (bhitk), to provide regular homography measurements, potentially further improving MOT performance while requiring only a camera-equipped device.
Furthermore, these contributions serve to afford robustness in situations where targets leave the ground plane; enabling tracking algorithms the best of both worlds: exploiting ground plane information when it is reliable, and relying solely on bounding boxes in the image plane otherwise, which is especially useful in applications like sports player tracking, dancing and gymnastic performances.
@@ 2 Related Work
IMM-JHSE uses only a single initial homography estimate as in (ucmc). It is important to note that the homography transformation only provides a plane-to-plane mapping. This differentiates it from fully-3D MOT methods which use regular 3D measurements obtained with LiDAR sensors (point clouds) (Jiang2024ATracking; Osep2018CombinedScenes; Zhang2024LEGO:Clouds; Wu2022CasA:Clouds; Nagy2024RobMOT:PointCloud; Huang2024BiTrack:Data), RGB-D cameras (Jafari2014Real-timeCameras; Luber2010LearningData), 3D detections (Reich2021MonocularTracks; Weng20203DMetrics; Wu20223DAssociation; Wang2024YouTracking; Wang2023TowardsMOT; Wu2023VirtualDetection; Wang2024MCTrack:Driving), odometry (Gao2021GlowEnvironments) and stereo or binocular cameras (Ninh2024CollabMOTTracking; Cao2015RobustCamera; Ess2009RobustPlatform).
In addition, the methods in (Choi2013ACamera; Wojek2013MonocularScenes) use a Reversible Jump Markov Chain Monte Carlo particle filter with various feature representations, including skin colour, depth images, facial detection (Choi2013ACamera) and explicit occlusion reasoning (Wojek2013MonocularScenes). Furthermore, these methods consider different camera motion models than the one proposed here (which explicitly considers camera position and rotation) and assume that internal camera parameters are known. Finally, (Wojek2013MonocularScenes) further makes use of odometry information.
Furthermore, the method in (Wu2023HybridDevices) attempts to perform 3D tracking without any of these measurements. However, the authors take an iterative approach to camera motion estimation. In contrast, IMM-JHSE integrates camera motion and its dynamics into the state vector to estimate it jointly with track states. They also rely on detecting vanishing points and lines parallel to the horizon. In addition, the multi-mode filter they employ only considers the target state and needs to be reinitialised with previous measurements when a mode switch occurs. Finally, multiple hypothesis tracking (MHT) performs association, whereas IMM-JHSE uses a linear assignment algorithm.
@@ 2.2 Tracking by attention
Instead of the pre-defined anchor boxes used in several popular object detectors, Carion et al. (end_end_object) decode object positions from a set of learned position embeddings using a transformer decoder. Meinhardt et al. (trackformer) extend this approach to MOT by using previous detections as queries in subsequent frames, resulting in the Trackformer architecture. Whereas Trackformer works on a frame-by-frame basis, Global Tracking Transformers, as proposed by Zhou et al. (gtr), encode detections in a sliding temporal window of frames. A transformer decoder then associates new detections with these encodings. Having access to an entire measurement sequence, the deep data association module proposed by Pinto et al. (transformer_dda) eschews the transformer decoder and relies solely on an encoder to associate detections within the given sequence with one another. End-to-end attention-based methods have recently succeeded on MOT leaderboards (fusiontrack; Gao2024MeMOTR:Tracking; Gao2024MultiplePrediction). However, these methods are computationally expensive, especially considering they often train their own detectors in addition to training the attention-based mechanism in an end-to-end approach.
@@ 2.3 Image-based Tracking
Tracking the motion of objects from a video has been of interest as early as 1998 when Isard and Blake (condensation) developed the condensation algorithm for tracking objects represented by a curve or sets of curves with learned motion models. Their method propagates conditional density distributions similar to the approach of particle filtering. However, tracking multiple objects simultaneously with their method may become infeasible, especially for complex multi-modal posterior densities where large samples or particles may be required to obtain accurate estimates. Besides, multi-object tracking requires solving the association problem – deciding which observations belong to which tracked object – which becomes increasingly difficult when the tracked objects are near one another, partially or fully occluding each other. Multiple object tracking (MOT) methods in the literature focus more on the association problem than on detection or individual target tracking.
@@
Given a set of object detections, Li et al. (Zhang2008GlobalFlows) model the association problem with a cost-flow network, with the solution obtained by a minimum-cost flow algorithm. They also explicitly model occluded observations, which is achieved by simply comparing differences in detection locations and scale to predetermined thresholds. While the approach is reported to be real-time and yields favourable results on the Context-Aware Vision using Image-based Active Recognition (CAVIAR) (caviar) dataset, object motion is not explicitly modelled. As a result, the method may perform poorly when object motion is highly dynamic. Berclaz et al. (Berclaz2011MultipleOptimization) similarly model object trajectories with a flow model consisting of a discretised 2D spatial grid (representing the image area) at each time step. They use the K-shortest path algorithm to solve their linear programming formulation of the association problem. Motion is modelled by only allowing an occupied node at the previous time step to transition to a specific neighbourhood around that node in the current time step. As with (Zhang2008GlobalFlows), this is an unsatisfactory solution in situations where objects are highly dynamic.
@@
With the increasing reliability and popularity of several object detectors (yolo; yolo9000; yolov3; yolov4; yolov9; Ge2021YOLOX:2021; end_end_object), modern methods have primarily made use of bounding box information. Bewley et al. (Bewley2016SimpleTracking) emphasise the importance of detection quality and use the Hungarian algorithm to perform data association based on the intersection-over-union (IoU) of detected bounding boxes with those predicted by Kalman-filtered tracks (specifically, predicted bounding boxes are constructed from the predicted bounding box centre x- and y-coordinates, height and aspect ratio). Their method has been influential in promoting the use of Kalman filters to estimate object motion distributions and/or the Hungarian algorithm (or other linear assignment algorithms) to perform association: these have become the status quo in (modelling_ambiguous; invisible; Wang2020TowardsTracking; Zhang2021FairMOT:Tracking; giao; Zhang2022ByteTrack:Box; botsort). Taking the minimalist approach of the original SORT algorithm one step further, Bergmann et al. (bells) rely purely on the regression head of their detector network (or “Tracktor”) to perform tracking. However, they show that performance improves with camera motion compensation and appearance-feature-based re-identification, suggesting that explicit motion models (and appearance information) can still be beneficial. Supporting this idea, Khurana et al. (invisible) use motion models that consider monocular depth estimates to maintain occluded tracks. SORT is improved upon in (Wojke2017SimpleMetric) by integrating appearance information, resulting in DeepSORT. They train a convolutional neural network (CNN) to discriminate between pedestrians and keep each track’s normalised appearance embeddings of the last 100 frames. Track association is performed by finding the minimum cosine distance of the current detected appearance embedding with the history of embeddings for each of the previous tracks. Additionally, they use motion information by obtaining the Mahalanobis distance between detections and predicted Kalman filter states, consisting of bounding box centre positions, aspect ratios, heights and their corresponding velocities. In SiMIlarity LEarning for Occlusion-Aware Multiple Object Tracking (SMILEtrack), Hsiang et al. (Wang2022SMILEtrack:Tracking) make use of a Siamese-based network which incorporates elements inspired by vision transformers (specifically, self-attention computed on image patches) to extract appearance features from detected bounding boxes. However, it does not achieve a convincing improvement in higher-order tracking accuracy (HOTA) (Luiten2021HOTA:Tracking) compared to ByteTrack, despite its superior computational speed. Leveraging multiple-camera traffic surveillance systems, another method explicitly models inter-vehicle occlusion using reconstructed-cuboid projections instead of relying on similarity learning (vehicle_occlusion). It is noted that the use of camera motion compensation has become ubiquitous, with applications in (bells; invisible; modelling_ambiguous; giao; Zhang2022ByteTrack:Box; botsort; Wang2022SMILEtrack:Tracking; interframe).
@@
Whereas previous methods separately perform detection and appearance embedding extraction, Wang et al. (Wang2020TowardsTracking) design a model which performs both of these simultaneously. They report that similar results can be achieved compared to state-of-the-art SDE (separate detection and embedding) methods at reduced computational cost. Zhang et al. (Zhang2021FairMOT:Tracking) show that such multi-task networks must be designed carefully since the tasks require different features and feature dimensions and that the anchors used for object detection can introduce ambiguity in the re-identification features. They introduce a carefully designed anchor-free model to simultaneously perform object detection and appearance embedding extraction, which outperforms the state-of-the-art methods in tracking metrics and frame rate.
@@
Shifting the focus from appearance embeddings, Zhang et al. (Zhang2022ByteTrack:Box) note that it is beneficial to perform association in two steps in their method called ByteTrack: the first step takes only high-confidence detections into account, while the remaining unmatched detections, as well as the lower-confidence detections, are associated in a second association step. They note that using the IoU is essential in the second association step since other features may become unreliable for low-confidence detections, which may be occluded and/or blurred. Still, re-identification or other features may be used in the first step. By altering the Kalman filter state vector, Aharon et al. (botsort) improve upon ByteTrack. Instead of tracking the aspect ratio of bounding boxes, they perform better by directly tracking bounding box widths and heights. Similar to (giao; Zhang2021FairMOT:Tracking; Wang2020TowardsTracking), an exponential moving average mechanism is used to update the appearance states of tracklets – the features of which are obtained by a re-identification network from the FastReID library (He2020FastReID:Re-identification). Specifically, the bags of tricks (Luo2019BagRe-Identification) stronger baseline with a ResNeSt50 (Zhang2020ResNeSt:Networks) backbone is used. Accordingly, they name their method BoT-SORT. Although tracking width and height may improve performance, Cao et al. (Cao2022Observation-CentricTracking) note that previous motion models are estimation-centric, allowing for rapid error accumulation over time, especially when no observations are available, and motion is non-linear. This implies that previous approaches are more sensitive to the noise in state estimation than the noise in measurements. They propose observation-centric SORT (OC-SORT), which incorporates a re-update step when an object is re-identified, during which the Kalman filter is updated with virtual detections generated based on the last-known and currently detected bounding box state. In addition, they use a novel association criterion that considers the change in motion direction that each new detection would induce, which is termed observation-centric momentum (OCM). Finally, Yang et al. (Yang2022HardSpace) buffers object bounding boxes to expand the matching space of subsequent detections in their Cascaded Buffered IoU (C-BIoU) tracker. A cascaded matching approach is used such that small buffers are used in the first association step, and larger buffers are used in the second step. Despite its simplicity, Their method performs superior to DeepSORT, SORT, ByteTrack and OC-SORT.